I am really stuck on this problem. I was hoping someone can help me solve it.
(4/3)(-6)(-3/4)
Im just not sure how to get the answer. The answer I came up with is 5, but Im not sure its right!

Answer by dabanfield(803)

(4/3)(-6)(-3/4) = ?
Let's do this in two steps:
First (4/3)*(-6) = -24/3
Now multiply -24/3 by -3/4
(-24/3)*(-3/4) = 72/12 = 6
Remember when you multiply two negatives (like -24 and -3) you get a positive.